OpenStudy (anonymous):

which is a factor of 12x^2-4x-8

OpenStudy (anonymous):

12x^2-4x-8=0 3x^2-x-2=0 3x^2-6x+x-2=0 3x(x-2)+1(x-2)=0 (x-2)(3x+1)=0

OpenStudy (anonymous):

thank you